OpenStudy (kebr1216):

Select all statements that are true about the linear equation. y=−x−1 A The graph of the equation is the set of points that are solutions to the equation. B The point (1, −2)(1, −2) lies on the graph of the equation. C The equation y+x=1y+x=1 has the same graph. D A coordinate pair on the graph of the equation is a solution to the equation. E The point (0, −1)(0, −1) lies on the graph of the equation.

OpenStudy (kebr1216):

oops i meant just y=−x−1

OpenStudy (danieldjpon3):

Oh ok

OpenStudy (danieldjpon3):

hold on brb ima help this other kid out real quick

OpenStudy (kebr1216):

ok ^_^

OpenStudy (danieldjpon3):

k im back

OpenStudy (kebr1216):

kk

OpenStudy (kebr1216):

its a multiple choice one

OpenStudy (danieldjpon3):

k

OpenStudy (danieldjpon3):

lol i don't get it.

OpenStudy (kebr1216):

i was thinking maby D and E but i feel like theirs more

OpenStudy (danieldjpon3):

hmm

OpenStudy (danieldjpon3):

@TheSmartOne

OpenStudy (danieldjpon3):

Can you help us out?

OpenStudy (danieldjpon3):

here brb

OpenStudy (danieldjpon3):

@@calculusxy

OpenStudy (danieldjpon3):

@calculusxy

OpenStudy (kebr1216):

@danieldjpon3 no one came T_T

OpenStudy (danieldjpon3):

ugh, just wait

OpenStudy (kebr1216):

kk ^_^

OpenStudy (danieldjpon3):

@agent0smith

OpenStudy (danieldjpon3):

we need help

OpenStudy (danieldjpon3):

ty for coming @agent0smith

OpenStudy (agent0smith):

Select all statements that are true about the linear equation. y=−x−1 A The graph of the equation is the set of points that are solutions to the equation. B The point (1, −2)(1, −2) lies on the graph of the equation. C The equation y+x=1y+x=1 has the same graph. D A coordinate pair on the graph of the equation is a solution to the equation. E The point (0, −1)(0, −1) lies on the graph of the equation. A is true B plug in x=1 and y=-2 into the equation and find out if that's true C simplify it and see if it's the same D true E Do the same thing as B

OpenStudy (kebr1216):

ok thx